# via Object-Centric Visual Representation Learning

### v2

@@ 3.2 UNO Architecture
An overview of UNO is in Fig. 2. First, we utilize a frozen pre-trained visual encoder to extract feature maps from the last layer of each video frame. Next, we apply Slot Attention [29] to decompose these feature maps into modular slots, effectively capturing both object and relation representations. To ensure spatio-temporal consistency, we introduce object temporal consistency learning, reinforcing stable slot features across the video. Finally, we propose a dynamic triplet prediction mechanism that associates relation slots with their corresponding object pairs. Integrated into our end-to-end framework, this mechanism minimizes redundancy in triplet prediction while enhancing model efficiency.
@@ 3.2.1 Visual Encoding
First, we employ a frozen, pre-trained vision model as the Visual Encoder, where the last-layer feature map encodes rich object cues [9, 46, 1, 43, 35], providing a spatial prior for learning object positions at varying granularities. Given a frame $I_{t}$, we extract its feature map $\mathbf{f}_{t}\in\mathbb{R}^{H_{\mathtt{enc}}\times W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$, where $H_{\mathtt{enc}}$ and $W_{\mathtt{enc}}$ denote the spatial dimensions, and $D_{\mathtt{enc}}$ represents the feature channel size.
Although pre-trained feature maps capture rich object information, they often entangle semantics, grouping similar objects [46, 9]. To address this, a decomposition module is essential for disentangling these features into distinct semantics, enabling precise object and relation predictions for triplet construction.
@@ 3.2.2 Object Decomposition
We employ Slot Attention [29], inspired by recent advances in object discovery [22, 60, 21], as a clustering mechanism to group semantically meaningful patches from $\mathbf{f}_{t}$ into predefined slots, where each slot corresponds to a distinct object region. Unlike previous studies [37, 21] that sample from a prior distribution, we initialize $N$ object slots (denoted as $\mathbf{s}$) as learnable tokens. We then decompose the feature map $\mathbf{f}_{t}$ of $t$-th frame into $N$ frame-wise object slot features $\mathbf{s}_{t}=\{\mathbf{s}_{t}^{1},\ldots,\mathbf{s}_{t}^{N}\}$, $\mathbf{s}_{t}\in\mathbb{R}^{N\times D{\mathtt{slot}}}$. This design encourages each slot to capture modular semantics [54, 18], enabling object-consistent decompositions through its unique competition mechanism that supports maintaining coherent slot representation across frames.
Formally, following the standard slot attention procedure [29], we employ three linear transformation heads to map the object slots $\mathbf{s}$ into Query $\mathbf{q}\in\mathbb{R}^{N\times D_{\mathtt{enc}}}$, while frame-wise feature maps $\mathbf{f}_{t}$ into Key $\mathbf{k}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$, and Value $\mathbf{v}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$. We iteratively calculate attention score and update slot representations via Gated Recurrent Unit (GRU) [5]. Mathematically, we formulate each iteration as:
$\tilde{\mathbf{a}}_{i,j}=\frac{e^{\mathbf{a}_{i,j}}}{\sum_{l=1}^{N}e^{\mathbf{a}_{l,j}}},\quad\text{where}\quad\mathbf{a}=\frac{1}{\sqrt{D}}\mathbf{q}\mathbf{k}^{\top}.$ (4)
@@
$\mathbf{s}_{t}=\text{GRU}(\texttt{inputs}=\mathbf{wv},\texttt{states}=\mathbf{s}_{t}),$
@@
Object Prediction. Since each object slot captures multi-level granularity information, using a lightweight prediction head is sufficient for object class, bounding box, and mask prediction. Thus, we employ three lightweight prediction heads, each consisting of a feed-forward network (FFN) with two linear layers–a functional layer followed by a task-specific layer for each output. The classification head outputs object classes $\hat{\mathbf{O}}_{t}=\text{FFN}_{\texttt{cls}}(\mathbf{s}_{t})$, while the box head predicts object coordinates $\hat{\mathbf{B}}_{t}=\text{FFN}_{\texttt{box}}(\mathbf{s}_{t})$. For mask prediction, a lightweight decoder with four transpose convolutions [30] upsamples the feature map $\mathbf{f}_{t}$ to the original frame size, resulting in $\mathbf{f^{\prime}}_{t}=\text{Dec}(\mathbf{f}_{t})$. The panoptic mask is then obtained by applying a matrix multiplication between the object slots $\mathbf{s}_{t}$ and the upsampled feature map $\mathbf{f^{\prime}}_{t}$ in the mask head, resulting in a binary mask of each object slot $\hat{\mathbf{m}}_{t}=\text{FFN}_{\texttt{mask}}(\mathbf{s}_{t}\cdot\mathbf{f^{\prime}}_{t})$.
Object Temporal Consistency Learning. Maintaining consistent spatio-temporal representations for object slots is crucial for tracking coherent object features over time. To achieve this, we employ object temporal consistency learning with a contrastive loss, inspired by prior works [23, 13]. Specifically, slots matching the same ground truth index across frames are treated as positive samples, while all others are negatives. Formally, given the $i$-th object slot $\mathbf{s}^{i}_{t}$ at frame $t$, we define $\mathbf{s}^{i}_{t-1}$ as the positive target and $\mathbf{s}^{j}_{t-1}$ as negatives ($j\neq i$). The right side of Fig. 2 illustrates the concept of object temporal consistency learning. The corresponding loss function is formulated as:
$\mathcal{L}_{\text{consistency}}=-\sum_{\mathbf{s}_{t-1}^{i}}\log\frac{e^{(\mathbf{s}_{t}^{i}\cdot\mathbf{s}_{t-1}^{i})}}{e^{(\mathbf{s}_{t}^{i}\cdot\mathbf{s}_{t-1}^{i})}+\sum_{\mathbf{s}_{t-1}^{j}}e^{(\mathbf{s}_{t}^{i}\cdot\mathbf{s}_{t-1}^{j})}}.$ (5)
@@ 3.2.3 Relation Decomposition
Existing VidSGG approaches [47] predict relations sequentially after object prediction, leading to high computational complexity and potential inaccuracy if objects are missed. In contrast, UNO predicts objects and relations simultaneously, enabling a parallelized execution that significantly reduces complexity. Using Eq. 4, we apply an enhanced slot attention mechanism to decompose the feature map $\mathbf{f}_{t}$ into relation slots, where each slot captures potential interaction regions. This mechanism enables relation slots to capture the entire spatial context of a frame, rather than being restricted to object pair intersections. This broader coverage significantly enhances relation prediction performance and supports our dynamic triplet mechanism (see Sec. 3.2.4). Specifically, we initialize $K$ relation slots (denoted as $\mathbf{z}$) as learnable tokens. Then, for the $t$-th frame, we obtain frame-wise relation slot features $\mathbf{z}_{t}=\{\mathbf{z}_{t}^{1},\ldots,\mathbf{z}_{t}^{K}\}$, where $\mathbf{z}_{t}\in\mathbb{R}^{K\times D{\mathtt{slot}}}$. Each frame-wise relation slot $\mathbf{z}_{t}$ is then passed through a classification FFN head to predict relations: $\hat{\mathbf{R}}_{t}=\text{FFN}_{\texttt{rel}}(\mathbf{z}_{t})$. Consequently, both relation slots $\mathbf{z}_{t}$ together with its relation class $\hat{\mathbf{R}}_{t}$ and object slots $\mathbf{s}_{t}$ are jointly extracted from $\mathbf{f}_{t}$ at time step $t$.
@@ 3.2.4 Dynamic Triplet Prediction
Existing methods either learn an adjacency matrix [50] to identify subject-object pairs or directly predict pairwise subject-object embeddings [47], followed by sequential relation prediction. In contrast, our approach introduces a dynamic triplet prediction mechanism that directly associates $N$ objects with $K$ relations, eliminating the need to construct an $N\times N$ object pair matrix. In theory, $K$ can be as large as $N^{2}$, since each object may interact with every other object. However, in practice, $K\ll N$ due to the inherent sparsity of real-world interactions—only a small subset of object pairs exhibit meaningful relationships. Detailed complexity analysis is provided in Appendix LABEL:app:efficiency. This reduces redundancy and duplication in triplet prediction while maintaining high performance. Fig. 3 illustrates the proposed module.
@@
$\mathbf{p}_{j}^{s}=\text{FFN}_{s}(\mathbf{z}_{t}^{j}),\quad\mathbf{p}_{j}^{o}=\text{FFN}_{o}(\mathbf{z}_{t}^{j})$ (6)
@@
$\hat{i}_{j}^{s}=\operatorname*{arg\,max}_{i\leq N}\>\text{sim}(\mathbf{p}_{j}^{s},\mathbf{s}^{i}_{t}),\>\hat{i}_{j}^{o}=\operatorname*{arg\,max}_{i^{\prime}\leq N}\>\text{sim}(\mathbf{p}_{j}^{o},\mathbf{s}^{i^{\prime}}_{t})$ (7)
@@
Triplet Duplication Reduction. Unlike prior work [47], which predicts triplets as a set and often results in duplicate detections, our method leverages slot attention with a built-in competitive mechanism to mitigate redundancy. This design ensures distinct object slots and relation slots focus on appropriate regions, producing unique object and relation predictions for triplets without requiring post-processing steps like Non-Max Suppression (NMS) [16].
@@ 3.3 Training Objectives
During training, we first perform Hungarian matching [11] between the predicted and ground-truth object boxes/masks to assign object slots, followed by supervision for detection and classification. The loss function for the DSGG task is defined as:
@@
where $\mathcal{L}_{\text{obj\_cls}}$ is the Cross Entropy (CE) loss, $\mathcal{L}_{\text{box}}$ is the $\ell_{1}$ loss, and $\mathcal{L}_{\text{GIoU}}$ is the GIoU loss [40].
@@
where $\mathcal{L}_{\text{mask}}$ is the CE loss, and $\mathcal{L}_{\text{dice}}$ is the Dice loss [44, 52].
@@
STTran [7] ICCV’21 ResNet-101 25.2 34.1 37.0 68.6 71.8 71.8 24.6 36.2 48.8 77.9 94.2 99.1
APT [25] CVPR’22 ResNet-101 26.3 36.1 38.3 69.4 73.8 73.8 25.7 37.9 50.1 78.5 95.1 99.2
STTran-TPI [51] ACM MM’22 ResNet-101 26.2 34.6 37.4 69.7 72.6 72.6
TR2 [48] ICRA’23 ResNet-101 26.8 35.5 38.3 70.9 73.8 73.8 27.8 39.2 50.0 83.1 96.6 99.9
VsCGG [31] ACM MM’23 ResNet-101 27.4 35.8 38.2 70.1 73.4 73.5 29.3 40.2 48.9 78.8 94.9 99.2
TEMPURA [32] CVPR’23 ResNet-101 28.1 33.4 34.9 68.8 71.5 71.5 29.8 38.1 46.4 80.4 94.2 99.4
DSG-DETR [12] WACV’23 ResNet-101 30.3 34.8 36.1 32.1 40.9 48.3
TPT [59] TMM’23 ResNet-101 32.0 39.6 51.5 85.6 97.4 99.9
@@
OED [47] CVPR’24 ResNet-50 33.5 40.9 48.9 73.0 76.1 76.1 35.3 44.0 51.8 83.3 95.3 99.2
UNO (Ours) ResNet-50 35.4 42.2 49.5 73.7 76.9 78.1 36.6 46.1 53.9 84.7 96.1 99.9
ViT-S/14 36.7 43.1 50.2 74.2 78.5 79.6 37.5 47.5 54.5 85.9 96.6 99.9
ViT-B/14 38.2 44.7 51.9 75.6 79.4 80.4 39.9 48.2 56.3 87.4 97.2 99.9
ViT-L/14 39.3 45.2 53.8 76.8 80.3 82.5 40.8 49.7 57.1 88.3 98.1 99.9
Table 3: Performance comparison with SOTA PVSG methods on PVSG dataset. The best results are in bold. Next, $\diamondsuit$ and $\heartsuit$ stand for the relation predictor [58]: 1D Convolution and Transformer Encoder, respectively.
@@
Multi-stage Method
Image Panoptic Segmentation + Tracking [4, 53]
$\diamondsuit$ PVSG [58] ResNet-50 3.88 5.24 6.71 2.55 3.29 5.36 10.06 14.99 18.13 8.98 12.21 15.47
$\heartsuit$ PVSG [58] ResNet-50 3.88 5.66 6.18 2.81 4.12 4.44 9.01 14.88 17.51 6.69 11.28 13.20
$\heartsuit$ MCL [34] ResNet-50 3.98 5.97 7.44 2.98 4.20 5.15 10.59 16.98 22.33 9.56 12.39 17.47
$\diamondsuit$ MCL [34] ResNet-50 4.51 6.08 7.76 3.56 4.38 5.86 11.43 17.30 22.85 9.57 13.13 17.48
Video Panoptic Segmentation [4, 24]
$\diamondsuit$ PVSG [58] ResNet-50 0.42 0.63 0.63 0.25 0.67 0.67 8.07 11.01 12.89 7.84 9.78 10.77
$\heartsuit$ PVSG [58] ResNet-50 0.42 0.73 1.05 0.61 0.76 0.92 6.50 9.64 12.26 5.75 8.25 9.51
$\heartsuit$ MCL [34] ResNet-50 0.63 1.05 1.05 0.83 0.76 0.76 6.71 10.27 13.42 6.94 8.68 12.09
$\diamondsuit$ MCL [34] ResNet-50 0.84 1.26 1.26 0.98 1.22 1.22 8.18 12.90 14.22 8.00 11.47 13.59

### v3

@@ 3.2 UNO Architecture
An overview of UNO is in Fig. 2. First, we utilize a frozen pre-trained visual encoder to extract feature maps from the last layer of each video frame. Next, we apply Slot Attention [locatello2020object] to decompose these feature maps into modular slots, effectively capturing both object and relation representations. These slots serve as shared latent features and are passed through a task-specific prediction head–either for bounding boxes or for panoptic masks. To ensure spatio-temporal consistency, we introduce object temporal consistency learning, reinforcing stable slot features across the video. Finally, we propose a dynamic triplet prediction mechanism that associates relation slots with their corresponding object pairs. Integrated into our end-to-end framework, this mechanism minimizes redundancy in triplet prediction while enhancing model efficiency.
@@ 3.2.1 Visual Encoding
First, we employ a frozen, pre-trained vision model as the Visual Encoder, where the last-layer feature map encodes rich object cues [ding2023betrayed, tian2023diffuse, caron2021emerging, simeoni2021localizing, oquab2023dinov2], providing a spatial prior for learning object positions at varying granularities. Given a frame $I_{t}$, we extract its feature map $\mathbf{f}_{t}\in\mathbb{R}^{H_{\mathtt{enc}}\times W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$, where $H_{\mathtt{enc}}$ and $W_{\mathtt{enc}}$ denote the spatial dimensions, and $D_{\mathtt{enc}}$ represents the feature channel size.
Although pre-trained feature maps capture rich object information, they often entangle semantics, grouping similar objects [tian2023diffuse, ding2023betrayed]. To address this, a decomposition module is essential for disentangling these features into distinct semantics, enabling precise object and relation predictions for triplet construction.
@@ 3.2.2 Object Decomposition
We employ Slot Attention [locatello2020object], inspired by recent advances in object discovery [scouter2021cvpr, zhou2022slot, actionslot2024], as a clustering mechanism to group semantically meaningful patches from $\mathbf{f}_{t}$ into predefined slots, where each slot corresponds to a distinct object region. Unlike previous studies [qian2023semantics, actionslot2024] that sample from a prior distribution, we initialize $N$ object slots (denoted as $\mathbf{s}$) as learnable tokens. We then decompose the feature map $\mathbf{f}_{t}$ of $t$-th frame into $N$ frame-wise object slot features $\mathbf{s}_{t}=\{\mathbf{s}_{t}^{1},\ldots,\mathbf{s}_{t}^{N}\}$, $\mathbf{s}_{t}\in\mathbb{R}^{N\times D{\mathtt{slot}}}$. This design encourages each slot to capture modular semantics [xu2022groupvit, DBLP:conf/iclr/JiaLH23], enabling object-consistent decompositions through its unique competition mechanism that supports maintaining coherent slot representation across frames.
Formally, following the standard slot attention procedure [locatello2020object], we employ three linear transformation heads to map the object slots $\mathbf{s}$ into Query $\mathbf{q}\in\mathbb{R}^{N\times D_{\mathtt{enc}}}$, while frame-wise feature maps $\mathbf{f}_{t}$ into Key $\mathbf{k}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$, and Value $\mathbf{v}\in\mathbb{R}^{H_{\mathtt{enc}}W_{\mathtt{enc}}\times D_{\mathtt{enc}}}$. We iteratively calculate attention score and update slot representations via Gated Recurrent Unit (GRU) [DBLP:conf/emnlp/ChoMGBBSB14]. Mathematically, we formulate each iteration as:
$\tilde{\mathbf{a}}_{i,j}=\frac{e^{\mathbf{a}_{i,j}}}{\sum_{l=1}^{N}e^{\mathbf{a}_{l,j}}},\qquad\text{where}\qquad\mathbf{a}=\frac{1}{\sqrt{D}}\mathbf{q}\mathbf{k}^{\top}.$ (4)
@@
$\mathbf{s}_{t}=\text{GRU}(\texttt{inputs}=\mathbf{wv},\texttt{states}=\mathbf{s}_{t})$
@@
Object Prediction. Since each object slot captures multi-level granularity information, using a lightweight prediction head is sufficient for object class, bounding box, and mask prediction. Thus, we employ three lightweight prediction heads, each consisting of a feed-forward network (FFN) with two linear layers–a functional layer followed by a task-specific layer for each output. The classification head outputs object classes $\hat{\mathbf{O}}_{t}=\text{FFN}_{\texttt{cls}}(\mathbf{s}_{t})$, while the box head predicts object coordinates $\hat{\mathbf{B}}_{t}=\text{FFN}_{\texttt{box}}(\mathbf{s}_{t})$. For mask prediction, a lightweight decoder with four transpose convolutions [DBLP:conf/cvpr/LongSD15] upsamples the feature map $\mathbf{f}_{t}$ to the original frame size, resulting in $\mathbf{f^{\prime}}_{t}=\text{Dec}(\mathbf{f}_{t})$. The panoptic mask is then obtained by applying a matrix multiplication between the object slots $\mathbf{s}_{t}$ and the upsampled feature map $\mathbf{f^{\prime}}_{t}$ in the mask head, resulting in a binary mask of each object slot $\hat{\mathbf{m}}_{t}=\text{FFN}_{\texttt{mask}}(\mathbf{s}_{t}\cdot\mathbf{f^{\prime}}_{t})$.
Object Temporal Consistency Learning. Maintaining consistent spatio-temporal representations for object slots is crucial for tracking coherent object features over time. Slot Attention provides a strong basis for object-centric video understanding. However, prior studies have noted that slot-based representations struggle to maintain temporal consistency across consecutive frames in a video [DBLP:conf/nips/ZadaianchukSM23]. To address this limitation, and inspired by prior works [DBLP:conf/cvpr/LiZPCCTL22, DBLP:journals/pami/FischerHPQCDY23], we incorporate object temporal consistency learning using a contrastive loss. Specifically, slots matching the same ground truth index across frames are treated as positive samples, while all others are negatives. Formally, given the $i$-th object slot $\mathbf{s}^{i}_{t}$ at frame $t$, we define $\mathbf{s}^{i}_{t-1}$ as the positive target and $\mathbf{s}^{j}_{t-1}$ as negatives ($j\neq i$). The right side of Fig. 2 illustrates the concept of object temporal consistency learning. The corresponding loss function is formulated as:
$\small\mathcal{L}_{\text{consistency}}=-\sum_{\mathbf{s}_{t-1}^{i}}\log\frac{e^{(\mathbf{s}_{t}^{i}\cdot\mathbf{s}_{t-1}^{i})}}{e^{(\mathbf{s}_{t}^{i}\cdot\mathbf{s}_{t-1}^{i})}+\sum_{\mathbf{s}_{t-1}^{j}}e^{(\mathbf{s}_{t}^{i}\cdot\mathbf{s}_{t-1}^{j})}}.$ (5)
@@ 3.2.3 Relation Decomposition
Existing VidSGG approaches [Wang_2024_CVPR] predict relations sequentially after object prediction, leading to high computational complexity and potential inaccuracy if objects are missed. In contrast, UNO predicts objects and relations simultaneously, enabling a parallelized execution that significantly reduces complexity. Using Eq. 4, we apply an enhanced slot attention mechanism to decompose the feature map $\mathbf{f}_{t}$ into relation slots, where each slot captures potential interaction regions. This mechanism enables relation slots to capture the entire spatial context of a frame, rather than being restricted to object pair intersections. This broader coverage significantly enhances relation prediction performance and supports our dynamic triplet mechanism (see Sec. 3.2.4). Specifically, we initialize $K$ relation slots (denoted as $\mathbf{z}$) as learnable tokens. Then, for the $t$-th frame, we obtain frame-wise relation slot features $\mathbf{z}_{t}=\{\mathbf{z}_{t}^{1},\ldots,\mathbf{z}_{t}^{K}\}$, where $\mathbf{z}_{t}\in\mathbb{R}^{K\times D{\mathtt{slot}}}$. Each frame-wise relation slot $\mathbf{z}_{t}$ is then passed through a classification FFN head to predict relations: $\hat{\mathbf{R}}_{t}=\text{FFN}_{\texttt{rel}}(\mathbf{z}_{t})$. Consequently, both relation slots $\mathbf{z}_{t}$ together with its relation class $\hat{\mathbf{R}}_{t}$ and object slots $\mathbf{s}_{t}$ are jointly extracted from $\mathbf{f}_{t}$ at time step $t$.
@@ 3.2.4 Dynamic Triplet Prediction
Existing methods either learn an adjacency matrix [panopticsg2024tpami] to identify subject-object pairs or directly predict pairwise subject-object embeddings [Wang_2024_CVPR], followed by sequential relation prediction. In contrast, our approach introduces a dynamic triplet prediction mechanism that directly associates $N$ objects with $K$ relations, eliminating the need to construct an $N\times N$ object pair matrix. In theory, $K$ can be as large as $N^{2}$, since each object may interact with every other object. However, in practice, $K\ll N$ due to the inherent sparsity of real-world interactions—only a small subset of object pairs exhibit meaningful relationships. This reduces redundancy and duplication in triplet prediction while maintaining high performance. Fig. 3 illustrates the proposed module.
@@
$\small\mathbf{p}_{j}^{s}=\text{FFN}_{s}(\mathbf{z}_{t}^{j}),\quad\mathbf{p}_{j}^{o}=\text{FFN}_{o}(\mathbf{z}_{t}^{j})$ (6)
@@
$\small\hat{i}_{j}^{s}=\operatorname*{arg\,max}_{i\leq N}\>\text{sim}(\mathbf{p}_{j}^{s},\mathbf{s}^{i}_{t}),\>\hat{i}_{j}^{o}=\operatorname*{arg\,max}_{i^{\prime}\leq N}\>\text{sim}(\mathbf{p}_{j}^{o},\mathbf{s}^{i^{\prime}}_{t})$ (7)
@@
Triplet Duplication Reduction. Unlike prior work [Wang_2024_CVPR], which predicts triplets as a set and often results in duplicate detections, our method leverages slot attention with a built-in competitive mechanism to mitigate redundancy. This design ensures distinct object slots and relation slots focus on appropriate regions, producing unique object and relation predictions for triplets without requiring post-processing steps like Non-Max Suppression (NMS) [DBLP:conf/cvpr/sangBS17].
@@ 3.3 Training Objectives
During training, we first perform Hungarian matching [DBLP:journals/jacm/EdmondsK72] between the predicted and ground-truth object boxes/masks to assign object slots, followed by supervision for detection and classification. The loss function for the DSGG task is defined as:
@@
where $\mathcal{L}_{\text{obj\_cls}}$ is the Cross Entropy (CE) loss, $\mathcal{L}_{\text{box}}$ is the $\ell_{1}$ loss, and $\mathcal{L}_{\text{GIoU}}$ is the GIoU loss [DBLP:conf/cvpr/RezatofighiTGS019].
@@
where $\mathcal{L}_{\text{mask}}$ is the CE loss, and $\mathcal{L}_{\text{dice}}$ is the Dice loss [DBLP:conf/miccai/SudreLVOC17, DBLP:conf/eccv/WangKSJL20].
@@
STTran [cong2021spatial] ICCV’21 ResNet-101 25.2 34.1 37.0 68.6 71.8 71.8 24.6 36.2 48.8 77.9 94.2 99.1
APT [li2022dynamic] CVPR’22 ResNet-101 26.3 36.1 38.3 69.4 73.8 73.8 25.7 37.9 50.1 78.5 95.1 99.2
STTran-TPI [wang2022dynamic] ACM MM’22 ResNet-101 26.2 34.6 37.4 69.7 72.6 72.6 - - - - - -
TR2 [wang2023cross] ICRA’23 ResNet-101 26.8 35.5 38.3 70.9 73.8 73.8 27.8 39.2 50.0 83.1 96.6 99.9
VsCGG [DBLP:conf/mm/LuCSLWH23] ACM MM’23 ResNet-101 27.4 35.8 38.2 70.1 73.4 73.5 29.3 40.2 48.9 78.8 94.9 99.2
TEMPURA [nag2023unbiased] CVPR’23 ResNet-101 28.1 33.4 34.9 68.8 71.5 71.5 29.8 38.1 46.4 80.4 94.2 99.4
DSG-DETR [feng2023exploiting] WACV’23 ResNet-101 30.3 34.8 36.1 - - - 32.1 40.9 48.3 - - -
TPT [zhang2023end] TMM’23 ResNet-101 - - - - - - 32.0 39.6 51.5 85.6 97.4 99.9
@@
OED [Wang_2024_CVPR] CVPR’24 ResNet-50 33.5 40.9 48.9 73.0 76.1 76.1 35.3 44.0 51.8 83.3 95.3 99.2
ResNet-50 35.4 42.2 49.5 73.7 76.9 78.1 36.6 46.1 53.9 84.7 96.1 99.9
ViT-S/14 36.7 43.1 50.2 74.2 78.5 79.6 37.5 47.5 54.5 85.9 96.6 100.0
ViT-B/14 38.2 44.7 51.9 75.6 79.4 80.4 39.9 48.2 56.3 87.4 97.2 100.0
UNO (Ours) ViT-L/14 39.3 45.2 53.8 76.8 80.3 82.5 40.8 49.7 57.1 88.3 98.1 100.0
Table 3: Performance comparison with SOTA PVSG methods on PVSG dataset. The best results are in bold. Next, $\diamondsuit$ and $\heartsuit$ stands for the relation predictor [yang2023pvsg]: 1D Convolution and Transformer Encoder, respectively.
@@
R@20↑ R@50↑ R@100↑ mR@20↑ mR@50↑ mR@100↑ R@20↑ R@50↑ R@100↑ mR@20↑ mR@50↑ mR@100↑
Multi-stage Method
Image Panoptic Segmentation + Tracking [cheng2021mask2former, wang2021different]
@@
[…]
Video Panoptic Segmentation [cheng2021mask2former, li2022videoknet]
[…]